# What Makes a Maze Look Like a Maze?

### v1

@@ 6 Discussion
Our framework demonstrates that by explicitly grounding conceptual schemas through large pretrained vision and language models, we can achieve better visual abstraction understanding in diverse real world instantiations. We find that LLMs generally possess well-structured knowledge of concepts, outputting comprehensive schemas that are accurately decomposed into more concrete symbols and their dependencies. The extracted schemas effective capture the concise essence of a concept, such as the answer to our initial question of what makes a maze look like a maze. However, as DSG relies on pre-trained LLMs to generate schemas and does not restrict the schemas to a set of specified symbols, it is possible that the schemas may contain harmful biases based on the concept given, or may include symbols that are difficult for VLMs to interpret.
In particular, we observe that current VLMs struggle to ground challenging schema components involving spatial constraints. DSG does not explicitly improve this capability or specify how to parameterize spatial configurations. While DSG can generate schema components that contain spatial relationships, and try to ground them by describing them in language, it remains limited by the VLMs’ abilities to conduct such spatial understanding. For example, let us consider a case where the extracted schema for the concept maze includes a symbol representing “an unobstructed path from entry to exit”. That such a path exists in the maze—typically only one and often one that is difficult to find—is a defining feature of mazes, yet is challenging for VLMs to correctly identify. VLMs’ lack of spatial understanding weakens DSG ability to interpret complex and spatially grounded components. Scaling DSG to better understand concepts that require precise layouts of schema components is difficult but crucial for future work.
@@
While we have made progress on understanding what makes a maze look like a maze, in proposing DSG as a framework for leveraging explicit schemas backed by powerful LLMs and VLMs, we are far from solving all reasoning across abstract concepts as humans do. There may be ways to improve grounding of individual components, as well as more efficient priors that can be incorporated into the system. We leave such exploration to future work.
@@ 7 Conclusion
We propose Deep Schema Grounding as a promising approach to understanding visual abstractions. DSG leverages schemas of concepts to decompose abstract concepts into subcomponents and model their dependencies. Our framework extracts explicitly structured schemas from large language models and hierarchically grounds them to images with vision-language models. On the Visual Abstractions Dataset, a visual question-answering benchmark composed of real-world images with diverse underlying abstract concepts, DSG demonstrates significant improvements compared to base VLMs. DSG is a step toward interpreting visual abstractions as humans do; more remains to be done in the challenge of visual abstraction reasoning.
@@ Acknowledgments
The appendix is organized as the following. In Appendix A, we present comparisons of Deep Schema Grounding to prior works with a graded accuracy metric, comparisons on free response generation, qualitative examples of DSG, and examples of hierarchical grounding. In Appendix B, we show examples of the Visual Abstractions Dataset, describe our annotation collection process, and release our dataset. In Appendix C, we provide our prompt to the LLM and all extracted schemas.
@@ A.1 Graded Accuracy Comparison
Due to variation in human responses on the Visual Abstractions Dataset—arising from natural errors in annotation or potential ambiguity in questions—we compare Deep Schema Grounding to prior works with a graded accuracy metric, based on degree of alignment to human judgements. Predicted answers each receive $\frac{K}{5}$ points, where $K$ is the number of human annotators out of $5$ that agree with the prediction. In Table 6, we report results with the graded accuracy metric over $5$ runs. DSG consistently outperforms previous works, showing an overall accuracy increase of $5.2$ percent point ($\uparrow 8.5\%$ relative improvement), and an accuracy boost in counting questions of $7.3$ percent point ($\uparrow 13.4\%$ relative improvement). We observe a similar trend in performance as with modal response accuracy reported in the main text—more improvement is seen in counting questions.
@@
GPT-4V (OpenAI, 2023) $0.611$ $0.546$ $0.648$ $0.603$
DSG (ours) $\mathbf{0.663}$ $\mathbf{0.619}$ $\mathbf{0.680}$ $\mathbf{0.663}$
@@ A.2 Comparison on Free Response Generation
We present comparisons between DSG and GPT-4V on free response generation in Table 7 and Table 8, with accuracy evaluated by BERTScores Zhang et al. (2020). We see that DSG outperforms GPT-4V across all questions and question types, and most concept categories.
@@
GPT-4V (OpenAI, 2023) $0.855$ $0.816$ $0.940$ $0.795$
DSG (ours) $\mathbf{0.865}$ $\mathbf{0.829}$ $\mathbf{0.952}$ $\mathbf{0.802}$
Table 8: DSG improves the performance of GPT-4V in free response generation across most categories of abstract concepts.
Strategic Scientific Social Domestic
GPT-4V (OpenAI, 2023) $0.813$$\pm 0.044$ $0.833$$\pm 0.080$ $0.903$$\pm 0.056$ $\mathbf{0.871}$$\pm 0.063$
DSG (ours) $\mathbf{0.828}$$\pm 0.055$ $\mathbf{0.845}$$\pm 0.066$ $\mathbf{0.917}$$\pm 0.052$ $0.870$$\pm 0.063$
A.3 Qualitative Examples
@@
A.4 Hierarchical Grounding Example
@@
A.5 Compute Resources
We use OpenAI’s API for GPT-4V (OpenAI, 2023), ViperGPT (Surís et al., 2023), and VisProg (Gupta & Kembhavi, 2023). Open-sourced models, LLaVA (Liu et al., 2024) and InstructBLIP (Dai et al., 2024), and the API calls of the aforementioned integrated LLMs with APIs, were run inference-only with $1$ A40 on an internal cluster.
Appendix B Visual Abstractions Dataset
B.1 Dataset Examples
In Figure 9, we provide visual question-answering examples of VAD across concept types and question types.
B.2 Annotation
We collect annotations for the Visual Abstractions Dataset through Prolific; $5$ human annotators labeled each question-image pair. Figure 8 shows our annotation interface, built from jsPsych. Each annotator was compensated at the rate of \$$\$12$ USD per hour, and no potential risk was incurred by the annotators.
@@ B.2 Annotation
B.3 Release
We release the Visual Abstractions Dataset here. We group all images by their corresponding abstract concepts. Each image is carefully curated and manually audited to be safe for release.
Figure 9: Examples of visual question-answering pairs in the Visual Abstractions Dataset.
@@ C.2 All Schemas
Here, we present all LLM-generated schemas for the Visual Abstractions Dataset.

### v2

@@ 6 Discussion
Our framework demonstrates that by explicitly grounding conceptual schemas through large pretrained vision and language models, we can achieve better visual abstraction understanding in diverse real-world instantiations. We find that LLMs generally possess well-structured knowledge of concepts, outputting comprehensive schemas that are accurately decomposed into more concrete symbols and their dependencies. The extracted schemas effectively capture the concise essence of a concept, such as the answer to our initial question of what makes a maze look like a maze. However, as DSG relies on pre-trained LLMs to generate schemas and does not restrict the schemas to a set of specified symbols, it is possible that the schemas may contain harmful biases based on the concept given, or may include symbols that are difficult for VLMs to interpret.
In particular, we observe that current VLMs struggle to ground challenging schema components involving spatial constraints. DSG does not explicitly improve this capability or specify how to parameterize spatial configurations. While DSG can generate schema components that contain spatial relationships, and try to ground them by describing them in language, it remains limited by the VLMs’ abilities to conduct such spatial understanding. For example, let us consider a case where the extracted schema for the concept maze includes a symbol representing “an unobstructed path from entry to exit”. That such a path exists in the maze—typically only one and often one that is difficult to find—is a defining feature of mazes, yet is challenging for VLMs to correctly identify. VLMs’ lack of spatial understanding weakens DSG’s ability to interpret complex and spatially grounded components. Scaling DSG to better understand concepts that require precise layouts of schema components is difficult but crucial for future work.
@@
While we have made progress on understanding what makes a maze look like a maze, in proposing DSG as a framework for leveraging explicit schemas backed by powerful LLMs and VLMs, we are far from solving all reasoning across abstract concepts as humans do. There may be ways to improve the grounding of individual components, as well as more efficient priors that can be incorporated into the system. We leave such exploration to future work.
@@ 7 Conclusion
We propose Deep Schema Grounding as a promising approach to understanding visual abstractions. DSG leverages schemas of concepts to decompose abstract concepts into subcomponents and model their dependencies. Our framework extracts explicitly structured schemas from large language models and hierarchically grounds them to images with vision-language models. On the Visual Abstractions Benchmark, a visual question-answering benchmark composed of real-world images with diverse underlying abstract concepts, DSG demonstrates significant improvements compared to base VLMs. DSG is a step toward interpreting visual abstractions as humans do; more remains to be done in the challenge of visual abstraction reasoning.
@@ Acknowledgments
The appendix is organized as the following. In Appendix A, we present comparisons of Deep Schema Grounding to prior works with a graded accuracy metric, comparisons on free response generation, results on single- v.s. multi-round reasoning, results with language-only baselines, ablations on schema complexity, qualitative examples of DSG, and examples of hierarchical grounding. In Appendix B, we show more examples of the Visual Abstractions Benchmark, explain our benchmark construction procedure, describe our annotation collection process, and release our dataset. In Appendix C, we provide our prompt to the LLM and all extracted schemas, as well as results from a human study evaluating the quality of the generated schemas.
@@ A.1 Graded Accuracy Comparison
Due to variation in human responses on the Visual Abstractions Benchmark—arising from natural errors in annotation or potential ambiguity in questions—we compare Deep Schema Grounding to prior works with a graded accuracy metric, based on degree of alignment to human judgements. Predicted answers each receive $\frac{K}{5}$ points, where $K$ is the number of human annotators out of $5$ that agree with the prediction. In Table 6, we report results with the graded accuracy metric over $5$ runs. DSG consistently outperforms previous works, showing an overall accuracy increase of $6.2$ percent points ($\uparrow 10\%$ relative improvement), and an accuracy boost in counting questions of $10.8$ percent points ($\uparrow 20.8\%$ relative improvement). We observe a similar trend in performance as with modal response accuracy reported in the main text—more improvement is seen in counting questions.
@@
GPT-4o (OpenAI, 2024) $0.622$ $0.519$ $0.655$ $0.632$
DSG (ours) $\mathbf{0.684}$ $\mathbf{0.627}$ $\mathbf{0.681}$ $\mathbf{0.708}$
@@ A.2 Comparison on Free Response Generation
We present comparisons between DSG and GPT-4o on free response generation in Table 7, with accuracy first evaluated by BERTScores Zhang et al. (2020). We see that DSG outperforms GPT-4o across all questions and question types.
@@
GPT-4o (OpenAI, 2024) $0.900$ $0.917$ $0.942$ $0.858$
DSG (ours) $\mathbf{0.908}$ $\mathbf{0.949}$ $\mathbf{0.946}$ $\mathbf{0.859}$
Additionally, we report free-response results by incorporating an LLM-based strategy for evaluating free-response answers (Kamalloo et al., 2023). In particular, we use zero-shot evaluation via LLM prompting, which includes the ground truth and predicted answer in the prompt, to verify the correctness of the prediction. We use GPT-4 as the base LLM, and compare GPT-4o with DSG. In Table 8, we see that DSG outperforms GPT-4o overall, with particularly strong results on counting and open-ended questions, and shows stronger signal than with BERTScores. However, performance on binary questions is lower, which we hypothesize is due to some of the components being potentially misleading or ambiguous in binary contexts.
Table 8: Performance on free response generation with  Kamalloo et al. (2023)’s GPT-4-based metric for evaluating open-domain QA models.
All Counting Binary Open
GPT-4o $36.67$ $45.56$ $\mathbf{55.24}$ $14.58$
DSG (ours) $\mathbf{37.78}$ $\mathbf{53.33}$ $49.52$ $\mathbf{21.25}$
A.3 Results on Single-round vs. Multi-round Reasoning
The key idea behind DSG lies in schema decomposition (e.g., in the form of extracting components: layout, walls, and entry and exit) to guide the VLM to interpret the scene. DSG introduces an explicit guidance to the VLM system, by first generating a decomposition of the target concept, parsing the image, then answering the question. The specific method of processing these questions—whether in a single query or across multiple queries—is not fundamentally important. While in theory, these steps could be merged into a single CoT-style prompt (Wei et al., 2022), our following experiments show that multi-round execution with explicit grounding steps significantly improves performance in practice, at least for the current vision-language models.
Here, we compare results of (1) GPT-4o, (2) a single-round CoT baseline given schema components (e.g. the VLM is given components for reasoning in a single query, without using the explicit grounding steps), (3) a single-round CoT baseline given the full hierarchical schema (also in Table 3), and (4) our full multi-round DSG framework. All results are averaged over 5 runs. In Table 9, we see that DSG’s multi-round execution outperforms single-round, CoT-style baselines, likely due to its explicit guidance and grounding.
Table 9: Comparison of DSG with single- and multi-round grounding of the schema, with results averaged over $5$ runs.
All Counting Binary Open
GPT-4o $0.664$ $0.604$ $0.655$ $0.693$
Single-round with schema components $0.667$ $0.607$ $0.661$ $0.696$
Single-round with hierarchical schema $0.681$ $0.629$ $0.672$ $0.709$
Multi-round with hierarchical schema (DSG) $\mathbf{0.730}$ $\mathbf{0.704}$ $\mathbf{0.690}$ $\mathbf{0.776}$
A.4 Language-only Baselines
We additionally report results on language baselines. The first is a language-only baseline, where GPT-4o only has access to the question itself without input image or schema. The second is a language-with-schema baseline, where GPT-4o has access to the question and the DSG grounded schema (the schema is extracted from the image and provides context), but the final query itself does not have access to the image. We report results over $5$ runs, and present results in Table 10.
Table 10: Comparison with language-only baseline and baseline with DSG’s grounded schema, with results averaged over $5$ runs.
All Counting Binary Open
Language-only $0.397$ $0.300$ $0.510$ $0.335$
Language w/ DSG schema $\mathbf{0.522}$ $\mathbf{0.440}$ $\mathbf{0.530}$ $\mathbf{0.546}$
Notably, the language-only baseline performs worse than all open-source and closed-source VLMs, showing that the visual input is essential for achieving higher accuracy on the benchmark. We note that the language-only baseline outperforms integrated LLMs with API baselines, as these methods tend to fail in execution. With DSG’s grounded schema, the language model performs significantly better, highlighting the importance of the holistic context provided by the grounded schema. We see that in open-ended questions, the grounded schema significantly improves performance of the language-only baseline. We hypothesize that this is because the schema helps rule out implausible answer choices by leveraging the grounded components.
A.5 Schema Complexity Ablation
We highlight that there’s a tradeoff between encouraging more detailed components in a schema and avoiding both components that are less broadly applicable across all image instantiations as well as potential failures of current VLMs in handling such detailed concepts, especially more spatially-focused ones. As capabilities of VLMs increase, we believe DSG’s performance will scale accordingly, and we can increase the detail of the schema without making this tradeoff. In addition, the complexity of the generated schema is currently influenced by the number of components in the example prompt. As LLMs become less dependent on the form of the initial prompt, they can decide the schema without being influenced by the bias in the example.
To explore the impact of schema complexity and better understand the capabilities of current VLMs, we report ablations results varying the numbers of components in the example prompt: $3$, $5$, $7$, and $9$. The example prompts are based on the same abstract concept academia expanded to different levels of detail.
With 3 components:
gen(concept=academia) =
gen(faculty | concept=academia)
gen(students | concept=academia)
gen(research-output | concept=academia, faculty, students)
With 5 components:
gen(concept=academia) =
gen(faculty | concept=academia)
gen(students | concept=academia)
gen(courses | concept=academia, faculty, students)
gen(research-output | concept=academia, faculty, students, courses)
gen(funding | concept=academia, research-output)
With 7 components:
gen(concept=academia) =
gen(faculty | concept=academia)
gen(students | concept=academia)
gen(departments | concept=academia, faculty, students)
gen(courses | concept=academia, departments, faculty, students)
gen(research-output | concept=academia, departments, faculty, students)
gen(funding | concept=academia, research-output, departments)
gen(infrastructure | concept=academia, funding, departments, courses)
With 9 components:
gen(concept=academia) =
gen(faculty | concept=academia)
gen(students | concept=academia)
gen(departments | concept=academia, faculty, students)
gen(courses | concept=academia, departments, faculty, students)
gen(research-output | concept=academia, departments, faculty, students)
gen(funding | concept=academia, research-output, departments)
gen(infrastructure | concept=academia, funding, departments, courses)
gen(administration | concept=academia, infrastructure, departments, faculty, students)
gen(community-engagement | concept=academia, research-output, administration, students, faculty)
Table 11: Ablation with varying number of components in the example prompt.
All Counting Binary Open
DSG w/ 3 comp. $\mathbf{0.730}$ $\mathbf{0.689}$ $\mathbf{0.700}$ $\mathbf{0.771}$
DSG w/ 5 comp. $0.691$ $0.622$ $0.671$ $0.733$
DSG w/ 7 comp. $0.685$ $0.667$ $0.662$ $0.713$
DSG w/ 9 comp. $0.676$ $0.700$ $0.657$ $0.683$
Table 12: Number of components in schema and the resulting average number of components in the generated schemas.
Generated comp. #
3 comp. $3.00$
5 comp. $4.66$
7 comp. $5.75$
9 comp. $7.00$
In Table 11, the results demonstrate that accuracy decreases as the number of components in the example prompt increases, likely due to current VLMs struggling with grounding more specific components. We note that all DSG variants still outperform the GPT-4o baseline. Interestingly, our findings also highlight a tendency for LLMs to generate shorter schemas, even when exposed to more detailed example prompts. In Table 12 we show the number components in the prompts and the mean number of components in the resulting schemas, averaged over all abstract concepts.
Notably, although the number of components in the example prompt does affect the average number of components in resulting schemas, this effect is less prominent and the resulting average plateaus. This phenomenon indicates that schemas may potentially converge to shorter lengths. And while current VLMs’ limitations constrain schema complexity, as these capabilities improve with development, we expect DSG’s performance to scale, enabling our framework to adopt more detailed schemas without sacrificing effectiveness.
A.6 Qualitative Examples
@@
A.7 Hierarchical Grounding Example
@@
A.8 Computing Resources
We use OpenAI’s API for GPT-4o (OpenAI, 2024), ViperGPT (Surís et al., 2023), and VisProg (Gupta & Kembhavi, 2023). Open-sourced models, LLaVA (Liu et al., 2024) and InstructBLIP (Dai et al., 2024), and the API calls of the aforementioned integrated LLMs with APIs, were run inference-only with $1$ A40 on an internal cluster.
Appendix B Visual Abstractions Benchmark
B.1 Benchmark Examples
In Figure 9, we provide visual question-answering examples of VAB across concept types and question types.
B.2 Benchmark Construction
We constructed the Visual Abstractions Benchmark inspired by areas where abstract concepts would occur in the real-world, including strategic games which can be played across instantiations, scientific concepts where many classroom science projects exist, social concepts with theory-of-mind examples, and domestic concepts with tasks that do not have predefined and exact configurations of end states. Images were manually curated from the Internet, primarily using Google Image Search. We searched for diverse real-world representations of each concept (e.g., tic-tac-toe on the beach). Images were chosen such that each image is (a) faithful to the given concept, (b) diverse compared to other chosen images, (c) real-world, and (d) high-quality; this process is difficult to procedurally replicate. We then wrote general questions for each concept, aiming for broad applicability and alignment with the images. These questions focus on reasoning grounded in the abstract concept rather than specific visual details of individual images. Once images and questions were curated, they were sent to human annotators via Prolific for labeling.
B.3 Annotation
We collect annotations for the Visual Abstractions Benchmark through Prolific; $5$ human annotators labeled each question-image pair. Figure 8 shows our annotation interface, built from jsPsych. Each annotator was compensated at the rate of \$$\$12$ USD per hour, and no potential risk was incurred by the annotators.
@@ B.3 Annotation
B.4 Release
We release the Visual Abstractions Benchmark here. We group all images by their corresponding abstract concepts. Each image is carefully curated and manually audited to be safe for release.
Figure 9: Examples of visual question-answering pairs in the Visual Abstractions Benchmark.
@@ C.2 All Schemas
Here, we present all LLM-generated schemas for the Visual Abstractions Benchmark.
@@
C.3 Human Study of Schema Quality
To evaluate the correctness and quality of the generated schemas, we conduct a human study using Prolific, with $20$ participants. Each participant was asked $3$ questions about each schema. The questions are as follow for an example concept maze:
Here you are given a concept, maze, and the core components underlying it: layout, walls, entry-exit. How well do these core components represent this concept on a scale of 1-7?
Here you are given a concept, maze, and the core components underlying it: layout, walls, entry-exit. How many components out of the ones listed accurately represent part of this concept?
Here you are given a concept, maze, and the core components underlying it: layout, walls, entry-exit. How many more components, if any, would you add to accurately represent this concept?
We collect $3$ responses for each of $12$ concepts with $20$ responses each, and yield a total of $720$ human annotated answers. In Table 13, we see that (a) participants rated schemas as reasonably representative, (b) $86.3\%$ of the components were considered accurate by participants, and (c) participants suggested adding $1.6$ components on average per schema. The concept that ranks highest for the first question is the concept cell with a rating of $6.15$, and lowest as the concept negotiating with $4.50$. The concept solar-system had the highest number of components to keep at $95\%$, while the concept negotiating is the lowest again at $61.7\%$. The concept tic-tac-toe was the most complete, with an average of $1$ concept that participants might add, while the concept solar-system had an average of $2.65$. Though there is still room for improvement, the schemas generated by the LLM are largely accurate and aligned with human expectations. Importantly, none of the schemas were judged as fully invalid, showing the robustness of DSG’s approach.
Table 13: Prolific study on quality of generated schemas, with results averaged over $20$ participants.
Representativeness (1-7) # comp. to keep # comp. to add
Participant responses $5.74$ $2.59$ $1.59$